# Merging and Disentangling Views in Visual Reinforcement Learning for Robotic Manipulation

### v1

@@ Abstract
Vision is well-known for its use in manipulation, especially using visual servoing. To make it robust, multiple cameras are needed to expand the field of view. That is computationally challenging. Merging multiple views and using Q-learning allows the design of more effective representations and optimization of sample efficiency. Such a solution might be expensive to deploy. To mitigate this, we introduce a Merge And Disentanglement (MAD) algorithm that efficiently merges views to increase sample efficiency while augmenting with single-view features to allow lightweight deployment and ensure robust policies. We demonstrate the efficiency and robustness of our approach using Meta-World and ManiSkill3.111Project Website and Code: here
Figure 1: Merge And Disentangle. We introduce a method that can merge multiple views during training to learn better representations, while simultaneously disentangling the view representations, such that the policy can function with any singular view input during deployment.
@@ 1 Introduction
Figure 2: Environment Setup. Our environment setup for robotic manipulation where we use three input camera views as inputs, namely First Person, Third Person A, and Third Person B across two visual RL benchmarks: (Left) Meta-World (Right) ManiSkill3. Extended visuals in Appendix D.
@@ 2 Related Works
Merging and Disentangling Features. Research in computer vision has extensively explored feature fusion  [45, 46, 47, 48] and feature disentanglement  [13, 49, 50] across different views and modalities. Inspired by these efforts, multiple works in visual RL explored feature merging  [51, 52, 44, 6] and feature disentanglement  [53, 17, 54]. However, policies trained on multiple feature inputs often suffer significant performance degradation when any feature inputs become unavailable  [55, 7]. Recently, Dunion and Albrecht [7] addressed this challenge in the multi-view RL setup for robotic manipulation. They introduced MVD, an actor-critic method that disentangles view representations into shared and private components through contrastive losses, enabling policies to maintain performance despite a reduction in input camera views. While MVD focuses exclusively on disentangling views, our approach focuses on both merging and disentangling views: we merge multi-view features to improve sample efficiency and we train on both merged multi-view features and singular-view features to ensure disentanglement in the case of a singular camera view input.
@@ 4.3 Augmentation Strategy
[…]
@@
[…]
@@ 5 Experiments
We benchmark our method and baselines on a total of 20 visual RL tasks, consisting of 15 Meta-World v2 tasks [9] and 5 ManiSkill3 tasks [10], focusing on tabletop robot manipulation tasks. Both environments are set up with three camera views, consisting of a first-person camera attached to the robot arm, and two third-person cameras with reasonable observability to the task at hand. See Figure 2 for visualizations of our camera setup in each environment. The full task list is defined in A.3, and detailed visuals of all tasks can be found in D. Through experimentation, our aim is to answer the following questions:
@@ 5.1 Empirical Results
Robustness. To quantify the sample efficiency of MAD, we train all methods on 20 visual RL tasks from Meta-World and ManiSkill3, using all three cameras as input, and evaluate the methods on all cameras views combined, and individual, reporting the combined, individual and average success rates in Figure 4. As our results indicate, MAD demonstrates superior performance to baselines on Meta-World by (30%) success rate over 15 tasks, and on ManiSkill3 by (36%) over 5 tasks. These improvements in success rates directly reflect the enhanced sample efficiency of MAD compared to baseline methods.
With a reduction in camera views, MAD achieves stronger robustness on the Third Person A and Third Person B singular views than all the baselines, including the Single Camera baseline that was trained solely on these third person views. This indicates that MAD was able to leverage better representations from All Cameras to achieve higher success rates on these third person views. On the First Person View, MAD outperforms most baselines and achieves similar success rates as the Single Camera and VIB baselines. Overall, MAD is able to achieve consistent success rates whether All Cameras are present, or when only singular camera view inputs are available, indicating a robustness to a reduction in camera views. Detailed graphs of all results can be found in Appendix C. Extended experiments on increasing views (5 views), and varying input view resolutions are provided in Appendix B.
@@ A.1 Hyper-parameters
Agent update frequency $2$
@@
Actor log stddev bounds $[-10,2]$
@@ B.2 Different Input Camera View Sizes
Appendix C Extended Results
C.1 Overall Robustness
@@
Overall Robustness. Success rate as a function of environment steps, averaged over 15 Meta-World and 5 ManiSkill3 visual RL tasks. (Top Left) Meta-World Medium. (Top Right) Meta-World Hard. (Bottom Left) Meta-World Very Hard. (Bottom Right) ManiSkill3. Methods are trained on all three camera views and evaluated on all and singular camera views. Mean and 95% CI over 5 random seeds. In reference to Figure 4 in the main text.
C.2 Ablations
@@
Ablations. Success rate as a function of environment steps, averaged over 5 Meta-World hard visual RL tasks. (Left) Alpha Ablations. (Right) Component Ablations. Methods are trained on all three camera views and evaluated on all and singular camera views. Mean and 95% CI over 5 random seeds. In reference to Figure 5 in the main text.
C.3 Multi-View Merging
Multi-view Merging. Success rate as a function of environment steps, averaged over 5 Meta-World hard visual RL tasks. Methods trained and evaluated on (Left) two camera views - First and Third A - (Right) three camera views. Mean and 95% CI over 5 random seeds. In reference to Figure 7 in the main text.
Appendix D Visuals
D.1 Meta-World

### v2

@@ Abstract
Vision is well-known for its use in manipulation, especially using visual servoing. Due to the 3D nature of the world, using multiple camera views and merging them creates better representations for Q-learning and in turn, trains more sample efficient policies. Nevertheless, these multi-view policies are sensitive to failing cameras and can be burdensome to deploy. To mitigate these issues, we introduce a Merge And Disentanglement (MAD) algorithm that efficiently merges views to increase sample efficiency while simultaneously disentangling views by augmenting multi-view feature inputs with single-view features. This produces robust policies and allows lightweight deployment. We demonstrate the efficiency and robustness of our approach using Meta-World and ManiSkill3.
Figure 1: Merge And Disentangle. We introduce a method that can merge multiple camera views during training to learn better representations, while simultaneously disentangling the camera view representations, such that the policy can function with any singular view input during deployment.
@@ 1 Introduction
Figure 2: Environment Setup. Our environment setup for robotic manipulation where we use three input camera views as inputs, namely First Person, Third Person A, and Third Person B across two visual RL benchmarks: (Left) Meta-World (Right) ManiSkill3. Extended visuals in Appendix E.
@@ 2 Related Works
Merging and Disentangling Features. Research in computer vision has extensively explored feature merging  [45, 46, 47, 48] and feature disentanglement  [13, 49, 50] across multiple camera views and sensor modalities. Inspired by these efforts, multiple works in visual RL have explored feature merging  [51, 52, 44, 6] and feature disentanglement  [53, 17, 54]. However, policies trained on multiple feature inputs often suffer significant performance degradation when any feature inputs become unavailable  [55, 7]. Multiple works have researched this problem of feature disentanglement for reinforcement learning. Hu et al. [56] explore the use of scaffolding through extra privileged features at training time, and disentangling them to achieve better test time performance than their unscaffolded counterparts. Skand et al. [57] introduce a multi sensor feature encoder for RL polices that disentangles features through dropout, yielding robust policies. Recently, Dunion and Albrecht [7] addressed feature disentanglement in the multi-camera view RL setup for robotic manipulation. They introduced MVD, an actor-critic method that disentangles view representations into shared and private components through contrastive losses, enabling policies to maintain performance despite a reduction in input camera views. While MVD focuses exclusively on disentangling views, our approach focuses on both merging and disentangling views: we merge multi-view features to improve sample efficiency and we train on both merged multi-view features and singular-view features to ensure disentanglement in the case of a singular camera view input.
@@ 4.3 Augmentation Strategy
$\mathcal{L}_{\pi_{\phi}}^{\textbf{MAD}}(\mathcal{D})=\alpha*\mathcal{L}_{\pi_{\phi}}^{\textbf{UnAug}}$ $(\mathcal{D})\;+\;(1-\alpha)*\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{\pi_{\phi}}^{\textbf{Aug}}(\mathcal{D},i)~~~{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\textrm{(actor)}}$ (1)
@@
$\mathcal{L}_{Q_{\theta}}^{\textbf{MAD}}(\mathcal{D})=\alpha*\mathcal{L}_{Q_{\theta}}^{\textbf{UnAug}}$ $(\mathcal{D})\;+\;(1-\alpha)*\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{Q_{\theta}}^{\textbf{Aug}}(\mathcal{D},i)~~~{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\textrm{(critic)}}$ (2)
@@ 5 Experiments
We benchmark our method and baselines on a total of 20 visual RL tasks, consisting of 15 Meta-World v2 tasks [9] and 5 ManiSkill3 tasks [10], focusing on tabletop robot manipulation tasks. Both environments are set up with three camera views, consisting of a first-person camera attached to the robot arm, and two third-person cameras with reasonable observability to the task at hand. See Figure 2 for visualizations of our camera setup in each environment. The full task list is defined in A.3, and detailed visuals of all tasks can be found in E. Through experimentation, our aim is to answer the following questions:
@@ 5.1 Empirical Results
Robustness. To quantify the sample efficiency of MAD, we train all methods on 20 visual RL tasks from Meta-World and ManiSkill3, using all three cameras as input, and evaluate the methods on all cameras views combined, and individual, reporting the combined, individual and average success rates in Figure 4. As our results indicate, MAD demonstrates superior performance to baselines on Meta-World by (30%) success rate over 15 tasks, and on ManiSkill3 by (36%) over 5 tasks. These improvements in success rates within the same number of environment steps as baseline methods directly reflect the enhanced sample efficiency of MAD.
With a reduction in camera views, MAD achieves stronger robustness on the Third Person A and Third Person B singular views than all the baselines, including the Single Camera baseline that was trained solely on these third person views. This indicates that MAD was able to leverage better representations from All Cameras to achieve higher success rates on these third person views. On the First Person View, MAD outperforms most baselines and achieves similar success rates as the Single Camera and VIB baselines. Overall, MAD is able to achieve consistent success rates whether All Cameras are present, or when only singular camera view inputs are available, indicating a robustness to a reduction in camera views. Detailed graphs of all results can be found in Appendix D. Extended experiments on increasing views (5 views), and varying input view resolutions are provided in Appendix B.
@@ A.1 Hyper-parameters
Agent update frequency 22
@@
Actor log stddev bounds [10,2][-10,2]
@@ B.2 Different Input Camera View Sizes
Appendix C Extended Analysis
C.1 Best Camera View for Learning
Is there a single camera view that consistently outperforms other camera views? Based on our experiments, there is a singular view that outperforms other singular views, and that being the First Person view. This observation aligns with findings from Hsu et al. [6]. Due to the first person camera’s attachment to the robot arm, and active perception system, it provides the best representations for robot learning. However, its attachment to the robot arm can be a double edged sword, as it can suffer from limited observability that prevents it from solving the task. We shed light on the PokeCube task from ManiSkill3 in Figure 11, where the First Person view isn’t sufficient to solve the task as it has limited observability. Nevertheless, MAD is able to leverage useful information from other third person views and achieve high success rate on All Cameras. In the case of a reduction in views, MAD can still maintain the highest possible success rate on the First Person view given its limited observability.
Figure 11: First Person Camera View Failure. Success rate as a function of environment steps on the ManiSkill3 PokeCube task. Methods are trained on all three views and evaluated an all and singular views separately. Mean and 95% CI over 5 random seeds.
Appendix D Extended Results
D.1 Overall Robustness
@@
Figure 12: Overall Robustness. Success rate as a function of environment steps, averaged over 15 Meta-World and 5 ManiSkill3 visual RL tasks. (Top Left) Meta-World Medium. (Top Right) Meta-World Hard. (Bottom Left) Meta-World Very Hard. (Bottom Right) ManiSkill3. Methods are trained on all three camera views and evaluated on all and singular camera views. Mean and 95% CI over 5 random seeds. In reference to Figure 4 in the main text.
D.2 Ablations
@@
Figure 13: Ablations. Success rate as a function of environment steps, averaged over 5 Meta-World hard visual RL tasks. (Left) Alpha Ablations. (Right) Component Ablations. Methods are trained on all three camera views and evaluated on all and singular camera views. Mean and 95% CI over 5 random seeds. In reference to Figure 5 in the main text.
D.3 Multi-View Merging
Figure 14: Multi-view Merging. Success rate as a function of environment steps, averaged over 5 Meta-World hard visual RL tasks. Methods trained and evaluated on (Left) two camera views - First and Third A - (Right) three camera views. Mean and 95% CI over 5 random seeds. In reference to Figure 7 in the main text.
Appendix E Visuals
E.1 Meta-World